# Rethinking Visual Layer Selection in Multimodal LLMs

### v1

@@ 4.3 Exp-II: Effect of Deep Layer
The last layer is not the optimal choice: As shown in Fig. 2, the last layer does not perform the best on any benchmark. Additional experimental results in Sec. 6.2 further validate this observation. The underlying reason lies in the CLIP model’s training mechanism, where supervision primarily focuses on aligning the final layer token with text embeddings. The token in the final layer is optimized by CLIP’s contrastive loss, making it highly specialized for the image-text matching task. However, this optimization process, driven by the attention mechanism, most significantly suppresses local details in the last layer.
Penultimate layer as the optimal choice: As shown in Fig. 2, the penultimate layer consistently achieves the best results. This is because it preserves more visual information while maintaining a text alignment capability second only to the final layer. This balance represents the optimal trade-off between visual richness and textual alignment.
Deep Layers are Essential for OCR Tasks
@@
As shown in Fig. 1, shallow layers provide negligible text information for the LLM. A clear boundary exists between the shallow and middle layers, with layer 12 marking the transition point. Layers before this point fail to contribute meaningfully to text processing, while a sharp performance gain occurs immediately afterward. This can be attributed to two essential requirements for OCR tasks:
@@
The visual representation has already aligned with the text space: Despite containing rich visual details, shallow layers lack intrinsic alignment with textual representations, limiting their usefulness for OCR tasks. Fig. 1(a) confirms this with notably low cosine similarity between shallow and deep layers, highlighting their distinct feature representations. This discrepancy poses a challenge for the connector, which can only align features originating from the deep (text) space or adjacent middle layers.
@@
5 Experiment: Effect of Data and Model Scale
To answer Q2: If the answer to Q1 is ‘yes’, does it still hold under different data and model scales? We extended our model on LLaVA-like structure but variants with different model sizes and using different training dataset scales.
@@ 5.1 Settings
More Training Data
Motivated by recent work [53], we aim to investigate the impact of training data scales. We adopt the same training procedure detailed in Sec. 4.1. In the first stage, we use the LLaVA 558k dataset [27]. For the second stage, we evaluate three dataset configurations: (1) LLaVA 665K, (2) Cambrian-1 737K [46], an expansion of the 665K dataset with additional OCR data, and (3) a custom 1M dataset that builds on the 737K dataset by incorporating data specifically curated for Counting and Spatial Relation tasks. The data for these tasks is sourced from CLEVR [17] and TallyQA [1]. The detailed composition of the dataset can be found in Appendix 9.
Different LLM Sizes
Building on the original 1.4B experiments, we extend our study to include MobileLLaMA [7] 2.7B and Vicuna v1.5 7B. These LLMs are selected due to their similar architectures, making them well-suited for investigating the impact of different LLM sizes.
LRS-based layer selection
@@
5.2 Exp-I: Effect of Data Scales
@@
Increasing data scale boosts the penultimate layer’s perception ability: After incorporating additional OCR and vision-centric data, the middle layers continue to outperform the penultimate layer, albeit with a decreasing gap. Since the visual encoder remains frozen during training, this suggests that compared to the middle layers, fine-grained information is less explicitly preserved in the deep layers. In other words, fine-grained details in the middle layers are more readily utilized by the LLM through the connector, whereas those in the deep layers are harder to extract, requiring larger amounts of data to activate the connector and enable the LLM to effectively capture these fine-grained features.
The penultimate layer does not always perform best, regardless of data scale: On the 665K dataset, layer 18 outperforms the penultimate layer by 3% on CVBench. After incorporating OCR training data and subsequently adding vision-centric data, it maintains its lead, though the margin narrows to 0.8% and 0.5%, respectively. on the 737k dataset, layer 18 occasionally outperforms penultimate layer, surpassing the penultimate layer by 0.8% in SEEDBench and 0.2% in GQA. Moreover, as shown in Appendix 8 and 9, results from just four layers across various subtasks already demonstrate that the penultimate layer does not consistently achieve the best performance, let alone when considering all 24 layers.
Increasing the data scale does not enhance the capability of shallow layers in OCR tasks: Layer 3 shows similar performance on OCR tasks before and after incorporating OCR training data, indicating that increasing task-specific data alone cannot overcome the inherent limitations of shallow representations. This is due to their natural misalignment with the textural feature space. In contrast, middle layers, though only partially aligned, still benefit from additional training.
Middle and deep layers exhibit minimal differences on hallucination task: Across all dataset sizes, shallow layers perform significantly worse than middle and deep layers. But, the performance difference between middle and deep layers is still negligible.
5.3 Exp-II: Effect of LLM Scales
In this section, we further validate that these findings hold across LLMs of different sizes. While conclusions from small models may not fully generalize to larger ones, the main findings still hold, as detailed below:
@@ 5.3 Exp-II: Effect of LLM Scales
The penultimate Layer remains the optimal choice in deep space regardless of LLM size: As shown in Fig. 4, the penultimate layer consistently outperforms the final layer across LLMs of 1.4B, 2.7B, and 7B, reinforcing our findings. This indicates that CLIP-ViT’s final-layer visual degradation, driven by its training paradigm, cannot be offset by a stronger LLM. Furthermore, increasing LLM size does not yield significant improvements in POPE performance, indicating that hallucination bottlenecks in MLLMs stem primarily from the quality of the visual representation.
Aligning with the prior results regardless of LLM size: As shown in Appendix 6 and 7, under the 2.7B model, the penultimate layer fails to achieve the best performance in half of the MME subtasks. For instance, Layer 18 performs better in Count, Position, and Existence tasks. Similarly, in SEEDBench, the penultimate layer underperforms in more than half of the subtasks, such as Spatial Relation and Instance Interaction. A similar trend is observed in the 7B model.
6 Experiment: Visual Feature Fusion
@@
In this section, we aim to address Q3: If Q1 and Q2 are affirmed, can we leverage the advantages of shallow and middle layers to design better MLLMs?
6.1 Settings
The equation below shows that the simplest visual feature fusion mechanism,
where denotes the set of selected layers, and each represents the feature representation extracted from layer with a dimension of . Here, is the number of visual tokens, and is the feature dimensionality of each token. The concatenation function merges these representations along the feature dimension, producing an output of size , where denotes the number of concatenated layers. The resulting is then fed into the Connector. Subsequently, we explore various fusion strategies to optimize feature integration.
6.2 Exp-I: Fusion Selection via LRS Analysis
The Layer-wise Representation Similarity (LRS) analysis reveals the potential of shallow and middle layers. Based on it, we selected representative layers from shallow, middle, and deep layers for fusion experiments, aiming to preliminarily explore the role of different representation spaces in fusion methods.
As shown in Tab. 3, we optimize our approach by selecting layers from different representation spaces for feature fusion. In the shallow layer, we select layers 1 and 3 as representatives. Layer 1, being the most chaotic, primarily captures early-stage visual features, while layer 3 is still in a chaotic state but performs relatively well. In the middle layer, we choose layers 18 and 17, as they achieve the first and second-best performance within this representation space. For the deep layer, we select layer 23, as it demonstrates the highest overall performance. The baseline configuration considers only layer 23 as the visual representation. We systematically explore different combinations of to analyze their impact:
1.
Incorporating a middle layer: .
2.
Incorporating middle, shallow layer: .
3.
Incorporating two middle layer: .
4.
Quantile-driven layer selection: .
5.
RLS-driven optimal combination: .
and our main findings can be summarized as follows:
Visual representation fusion improves performance: As demonstrated in Tab. 3, all fusion methods exhibit a substantial performance improvement over the baseline. The best fusion approach outperforms the baseline in 9 out of 10 benchmarks. These results suggest that a direct visual feature fusion strategy can effectively enhance performance across various task categories, including General, OCR, vision-centric, and Hallucination tasks.
On Layer Combinations: Compared to two-layer fusion (), three-layer fusion () results in more stable performance improvements. However, the inclusion of shallow features diminishes the effectiveness of fusion strategies in OCR tasks. Feature layer selection guided by LRS analysis consistently outperforms selection methods based on quantiles (). As demonstrated in Sec. 4.3, the representation space of shallow layers has a minimal impact on OCR tasks. Therefore, incorporating shallow layers into feature fusion may negatively affect performance.
@@ 6.3 Exp-II: Fusion Method Comparison
Under the same experimental conditions, we select different combinations of visual representations based on LRS analysis. Despite its simplicity, this approach consistently surpasses existing fusion strategies across multiple benchmarks, underscoring the significance of our findings.
We evaluate four state-of-the-art fusion methods, including three from DenseConnector (DC) [49] and one from MMFuser [4]. As shown in Tab. 3, our method outperforms STI, SCI, and DCI on 8, 7, and 7 benchmarks, respectively. Compared to MMFuser [4], our approach demonstrates superior performance on 8 benchmarks. These results highlight the significant potential of visual feature fusion strategies in enhancing MLLMs and offering guidance for developing future fusion strategies.
@@ 7 Conclusion
This study presents a comprehensive layer-wise analysis, revealing that shallow and middle representation spaces can surpass the performance of deep layers. Evaluations across diverse data and model scales further substantiate this finding. Furthermore, we introduce a straightforward yet highly effective fusion strategy for visual feature integration, delivering substantial improvements over the baseline. Our findings offer a foundation for advancing future research in fusion methodologies.
Supplementary Material
We provide some additional information as a supplementary material. This material is divided into three sections:
Detailed analysis of visual representations
Experiment Details
Guidelines on Alignment
8 Visual Representations
In this section, we explore the variations in visual representation spaces across four tasks. As illustrated in Figure 5, the partitioning of visual representations is minimally influenced by the nature of the tasks. In other words, the shallow, middle, and deep representation spaces exhibit remarkable stability, maintaining consistent structures across various tasks.
Figure 6 further highlights the distinct characteristics of different representation spaces through the results on TextVQA and OCRBench. These results clearly demonstrate that shallow layers are ineffective for OCR tasks, with performance progressively improving in the middle space and being adequately addressed only in the deep space.
Another example, as shown in Table 4, when the penultimate layer visual representations are replaced with those from other layers of the visual encoder, layers 20 to 24, belonging to the deep visual representation space, show no signs of catastrophic performance degradation.
@@ 8 Visual Representations
| LLaVA [29] | 158K |
| + ShareGPT [42] | 40K |
| + VQAv2 [11] | 83K |
| + GQA [14] | 72K |
| + OKVQA [34] | 9K |
| + OCRVQA [36] | 80K |
| + A-OKVQA [41] | 66K |
| + TextCaps [43] | 22K |
| + RefCOCO [19, 33] | 48K |
| + VG [21] | 86K |
| LLaVA-1.5 [26] | 665K |
| + AI2D [20] | 16K |
| + DocVQA [35] | 15K |
| + DVQA [18] | 13K |
| Cambrian-737k [46] | 737K |
| + CLEVR [17] | 215k |
| + TallyQA [1] | 77K |
@@
9 Experiment Details
9.1 Visualization of Fusion Performance
As shown in Fig. 7, our fusion strategy achieves highly competitive performance.
9.2 Composition of Three Scale Datasets
The following datasets are incorporated to enhance the model’s capabilities across multiple multimodal tasks:
AI2D (Allen Institute for AI Diagram Dataset) [20] AI2D is designed for visual reasoning and diagram understanding, featuring annotated diagrams with textual descriptions and Q&A pairs. It is particularly useful for multimodal reasoning and visual question answering (VQA) tasks.
DocVQA (Document Visual Question Answering) [35] DocVQA focuses on visual question answering over document images, where questions pertain to scanned documents, OCR-recognized text, and textual reasoning. This dataset is valuable for document comprehension, text recognition, and multimodal reasoning.
DVQA (Diagrammatic Visual Question Answering) [18] DVQA is designed for visual question answering over diagrams and charts, covering questions related to bar charts, pie charts, and scientific illustrations. It evaluates the model’s ability to read structured visual information and perform reasoning based on graphical representations.
CLEVR (Compositional Language and Elementary Visual Reasoning) [17] CLEVR is a synthetic dataset for visual reasoning, containing 3D-rendered scenes with structured questions that require reasoning based on attributes, object relationships, and compositional logic. It is widely used to assess a model’s capability in compositional and multi-step reasoning.
TallyQA [1] TallyQA is a dataset specifically designed for object counting tasks, where questions require the model to accurately count objects in an image. It evaluates the model’s ability to attend to relevant objects, integrate global and local information, and perform numerical reasoning.
9.3 Evaluation Metrics

### v2

@@ 4.1 Deep-to-Deep Layer Comparison
The underlying reason might lie in the CLIP model’s training mechanism, where supervision primarily focuses on aligning the final layer token with text embeddings. The token in the final layer is optimized by CLIP’s contrastive loss, making it highly specialized for the image-text matching task. However, this optimization process can significantly suppress local image features in the final layer, which can harm performance when finer-grained details are required.
Penultimate layer as the optimal choice.
As shown in Fig. 2, the penultimate layer consistently achieves the best performance across tasks. Notably, it outperforms other deep layers on General, OCR, and vision-centric tasks. This superiority stems from its ability to retain rich visual information while maintaining strong text alignment, which is second only to the final layer. Such a balance offers an optimal trade-off between visual expressiveness and semantic alignment, making it particularly well-suited for multimodal tasks.
4.2 Deep-to-Shallower Layer Comparison
Afterwards, we investigate the effects of using shallow and middle layers in MLLMs. We highlight the following key observations.
Deep layers are essential for OCR.
@@ 4.2 Deep-to-Shallower Layer Comparison
As shown in Fig. 1, shallow layers provide negligible text information for the LLM. A clear boundary exists between the shallow and middle layers, with layer 12 marking the transition point. Layers before this point fail to contribute meaningfully to text processing, while a sharp performance gain occurs immediately afterward. This might be attributed to two essential requirements for OCR tasks:
1.
@@
2.
Well textually aligned visual features: Despite containing rich visual details, shallow layers lack intrinsic alignment with textual representations, limiting their usefulness for OCR tasks. Fig. 1(a) confirms this with notably low cosine similarity between shallow and deep layers. This discrepancy poses a challenge for the connector, which can only align features originating from the deep (text) space or adjacent middle layers.
Limited impact of representation quality on cognitive tasks.
Cognitive tasks, such as “Code Reasoning” and “Numerical Calculation”, require both perception and high-level reasoning capabilities. Interestingly, we observe that even the shallow layers, which generally yield lower quality visual representations can rival or even outperform deeper layers in tasks under the MME-Cognitive. layer 3 achieves superior performance on “Code Reasoning”, “Numerical Calculation’, and “Text Recognition” compared to both middle and deep layers. (see Tab.7 and Tab.9 in Appendix) These findings suggest that, for cognitive tasks, visual feature quality is not the primary limiting factor.
Potential of middle layers
We first conduct an investigation into the middle representation spaces. The performance of these two spaces, as shown in Tab. 1, several key insights emerge from the results:
(1) The middle layer has the potential to perform best: Although the middle layer’s information has not been fully processed, it still achieves the best performance on one-third of the benchmarks. Specifically, compared to the penultimate layer, layer 14 achieves a 20-point higher score on MME-Cognitive, layer 18 outperforms by 3% on CVBench, layer 17 surpasses by 1.4% on RefCOCO, and layer 20 exceeds by 0.2% on GQA.
(2) The middle layers generally perform better on vision-centric tasks: Fig. 3(a) illustrates the performance of different layers across subtasks in the MME dataset, showing that position and existence tasks benefit more from middle layer representations. As depicted in Fig. 3(b), the penultimate layer achieves top performance in only three out of eleven subtasks, whereas the shallow and middle layers yield optimal results in seven. Similarly, in Fig. 3(c), one-third of the best performing results, such as those in spatial relations, physical relations, and cross fine-grained perception originate from the shallow and middle layers. A comparable trend is observed in SEEDBench (Fig. 3(d)), where middle layers produce optimal results in nearly half of the subtasks, including Instance Attribute, Instance Location, Instance Interaction, and Text Recognition.
The hallucination problem is more pronounced in shallow layers but is effectively mitigated in the middle layers.
As shown in Tab. 1, POPE results indicate that hallucination issues are most prominent in the shallow representation space, with minimal variation between the middle and deep layers. Notably, in the middle representation space, half of the layers outperform the penultimate layer on the POPE. This phenomenon likely stems from the fact that the challenge of this task lies more in visual perception than in semantic comprehension. In Sec.5.3, we provide a detailed analysis showing that further experiments with larger LLMs consistently support this finding.
@@
5 Effect of Data and Model Scale
To further assess the generality of our findings, we extend our experiments to larger model scales and training datasets in this section and analyze the resulting performance trends.
@@ 5.1 Settings
More training Data.
We investigate the impact of training data scales using the same training procedure detailed in Sec. 3. In the first stage, we use the LLaVA 558k dataset Liu et al. (2023b). For the second stage, we evaluate three dataset configurations: (1) LLaVA 665K, (2) Cambrian-1 737K Tong et al. (2024a), an expansion of the 665K dataset with additional OCR data, and (3) a custom 1M dataset that builds on the 737K dataset by incorporating data specifically curated for vision-centric tasks. The dataset composition can be found in Appendix B.
Scaling LLM sizes.
Building on the original 1.4B experiments, we extend our study to include MobileLLaMA Chu et al. (2024) 2.7B and Vicuna v1.5 7B. These LLMs are selected due to their similar architectures, making them well-suited for investigating the impact of different LLM sizes.
@@
5.2 Deep-to-Deep Layer Comparison
@@
The penultimate layer remains the optimal choice in deep space regardless of LLM size
@@
As shown in Fig. 4, the penultimate layer consistently outperforms the final layer across LLMs of 1.4B, 2.7B, and 7B, reinforcing our findings. This indicates that CLIP-ViT’s final-layer visual degradation, driven by its training paradigm, cannot be offset by a stronger LLM.
Furthermore, increasing LLM size does not yield significant improvements in POPE performance, indicating that hallucination bottlenecks in MLLMs stem primarily from the quality of the visual representation. In contrast, vision-centric tasks benefit more from scaling LLM size, indicating that even for tasks grounded in visual understanding, strong perception alone does not suffice, as robust reasoning capabilities remain essential.
The more data and the larger the model, the more the deep layers benefit.
As both the training data scale and LLM size increase, a clear trend emerges in Fig. 5: the proportion of subtasks where the penultimate layer achieves the best performance consistently grows.
Since the visual encoder remains frozen during training, this suggests that compared to the middle layers, fine-grained information is less explicitly preserved in the deep layers. In other words, fine-grained details in the middle layers are more readily utilized by the LLM, whereas those in the deep layers are harder to extract, requiring larger amounts of data to activate the LLM to effectively capture these fine-grained features.
5.3 Deep-to-Shallower Layer Comparison
We observe that while conclusions from small models may not fully generalize to larger ones, the main findings still hold, as detailed below:
The potential of shallower layers persists across model and data scales
The consistency of the previous conclusion is validated under larger training data and increased LLM size. As detailed in Appendix 9 and 10, under the 2.7B model, the penultimate layer fails to outperform shallower layers on several MME subtasks, such as Count, Position, and Existence. Similar patterns emerge in SEEDBench and persist with the 7B model, where shallower layers (e.g., Layer 18) achieve better results on tasks like Spatial Relation. On the 665K dataset, layer 18 outperforms the penultimate layer by up to 3% on CVBench and maintains a slight advantage on OCR and vision-centric tasks. This trend persists on the 737K dataset, where layer 18 continues to lead on SEEDBench and GQA. Although the performance gap narrows on the 1M dataset, the middle layer still surpasses the deep layer on both OCRBench and CVBench.
Limited gains on OCR tasks despite data and LLM scaling.
Layer 3 exhibits comparable performance on OCR tasks both before and after incorporating OCR-specific training data, and across models of different scales (2.7B vs. 7B). This suggests that increasing task-specific data alone cannot overcome the inherent limitations of shallow representations. The lack of improvement can be attributed to their poor alignment with the textural feature space required for OCR understanding. In contrast, the middle layers, though only partially aligned, still exhibit performance gains under additional supervision.
@@ 5.3 Deep-to-Shallower Layer Comparison
6 Visual Feature Fusion
Building on the above-mentioned findings that deeper layers are not universally optimal and that shallower layers offer valuable complementary information, we explore the most effective way to enhance visual representation by combining visual features from multiple layers. To be specific, we employ a simple fusion strategy to merge features from different layers and conduct a preliminary study on various layer combinations, aiming to highlight the potential benefits of layer fusion.
6.1 Method
The equation below shows the simplest visual feature fusion mechanism,
where denotes the set of selected layers, and each represents the feature representation extracted from layer with a dimension of . Here, is the number of visual tokens, and is the feature dimensionality of each token. The concatenation function merges these representations along the feature dimension, producing an output of size , where denotes the number of concatenated layers. The resulting is then fed into the Connector. Subsequently, we explore various layer combinations.
6.2 Exp-I: Ablation of Fusion Layer Selection
We select representative layers from shallow, middle, and deep layers for the fusion ablation study to preliminarily explore the effect of different representation spaces in fusion methods. We systematically construct different layer combinations to analyze their impact on fusion performance.
Multiple stages bring generalization:
Fig. 6 illustrates six different configurations, ranging from using only the end stage to incorporating all stages. Compared to two-layer fusion () and three-layer fusion covering two stages (), three-layer fusion () that spans all stages ( and Ours {23, 3, 18}) leads to more consistent performance improvements. Notably, performs worse than Ours {23, 3, 18} on OCR tasks. This is likely because the first layer feature is too raw, making them less suitable for extracting the low-level visual cues required in OCR. By incorporating more stable representations from multiple stages, models can achieve the most robust performance across tasks.
@@ 6.3 Exp-II: Fusion Method Comparison
To investigate the impact of different layer fusion strategies, we compare our selected layers {23, 3, 18} and a simple concatenation strategy, against other carefully designed fusion methods.
Less is more:
We compare two state-of-the-art ad-hoc methods, DenseConnector (DC) and MMFuser with our method. As shown in Tab. 3, our method outperforms DC and MMFuser on 7 and 8 across 10 widely used benchmarks, respectively. These results suggest that complex fusion strategies may be unnecessary, as the simplest concatenation already meets the performance requirements.
@@ 7 Conclusion
This study presents a comprehensive layer-wise analysis, revealing that shallow and middle representation spaces can surpass the performance of deep layers. Evaluations across diverse data and model scales further substantiate our findings. Furthermore, we introduce a straightforward yet highly effective fusion strategy for visual feature integration, delivering substantial improvements over the baseline. Our findings offer a foundation for advancing future research in fusion methodologies.
Limitations
Due to the high computational cost of layer-wise analysis, we adopt a linear probing-inspired strategy: most experiments are conducted on the 1.4B model, with selective validation on the 2.7B and 7B variants. However, our study does not extend to larger-scale LLMs. In terms of visual encoders, we focus exclusively on CLIP-ViT-L/14, given its widespread adoption, and leave the exploration of alternative backbones to future work.
Moreover, vision–language fusion strategies can be broadly classified into internal and external methods. Our analysis focuses only on external fusion approaches and does not directly compare them with internal alternatives. We ensure consistent experimental conditions across all settings, which enables a fair assessment of visual representation quality. However, our current design does not explore how different connector architectures may affect performance. We regard this as a valuable direction for future research.
Acknowledgement
We are grateful to EIT and IDT High Performance Computing Center for providing computational resources that greatly facilitated this research. This work was financially supported by the 2035 Key Research and Development Program of Ningbo City (Grant No. 2025Z034), the Pioneer and Leading Goose R&D Pr ogram of Zhejiang (Grant No. 2024C01110), and the National Natural Science Foundation of China (Grant No. 62472385).
Appendix
We provide some additional information as supplementary material. This material is divided into four sections:
A detailed analysis and formal computation process of visual representations is provided in Appendix A.
Experimental details covering dataset composition, evaluation protocols, visualization of fusion performance, and analyses on model factors such as LLM choice, data scale, layer selection, and feature fusion are provided in Appendix B.
We provide additional details of our reasoning and rebuttal explanations in Appendix C.
Ethics statement is provided in Appendix D.
Appendix A Visual Representations
A.1 Formal Computation of LRS
To examine the behavioral patterns of different visual layers, we introduce Layer-wise Representation Similarity (LRS) to analyze the relationships between various visual layers.
Specifically, we compute the cosine similarity between hidden states of ViT layers, where values closer to indicate higher similarity. The formalized division process is detailed in the following equation:
We define the hidden state matrix of CLIP-ViT as:
We also explore the variations in visual representation spaces across four tasks. As illustrated in Figure 7, the partitioning of visual representations is minimally influenced by the nature of the tasks. In other words, the shallow, middle, and deep representation spaces exhibit remarkable stability, maintaining consistent structures across various tasks.
A.2 Replacing the Penultimate Layer
As shown in Table 4, when the penultimate-layer visual representations are replaced by those from other layers of the visual encoder without additional training, layers 20 to 24 (belonging to the deep representation space) do not suffer catastrophic performance degradation. Moreover, among the two layers closest to the penultimate one, layer 22 exhibits more stable performance than the final layer.
@@ A.2 Replacing the Penultimate Layer
| LLaVA Liu et al. (2024b) | 158K |
| + ShareGPT ShareGPT (2023) | 40K |
| + VQAv2 Goyal et al. (2017) | 83K |
| + GQA Hudson and Manning (2019b) | 72K |
| + OKVQA Marino et al. (2019) | 9K |
| + OCRVQA Mishra et al. (2019) | 80K |
| + A-OKVQA Schwenk et al. (2022) | 66K |
| + TextCaps Sidorov et al. (2020) | 22K |
| + RefCOCO Kazemzadeh et al. (2014); Mao et al. (2016) | 48K |
| + VG Krishna et al. (2017) | 86K |
| LLaVA-1.5 Liu et al. (2023a) | 665K |
| + AI2D Kembhavi et al. (2016) | 16K |
| + DocVQA Mathew et al. (2021) | 15K |
| + DVQA Kafle et al. (2018) | 13K |
| Cambrian-737k Tong et al. (2024a) | 737K |
| + CLEVR Johnson et al. (2017) | 215k |
| + TallyQA Acharya et al. (2019) | 77K |
@@
Appendix B Experiment Details
B.1 Composition of Three Scale Datasets
As shown in Table 5, the following datasets are incorporated to enhance the model’s capabilities across multiple multimodal tasks, with all datasets accessed and utilized strictly under their licenses.
AI2D (Allen Institute for AI Diagram Dataset) Kembhavi et al. (2016) AI2D is designed for visual reasoning and diagram understanding, featuring annotated diagrams with textual descriptions and Q&A pairs. It is particularly useful for multimodal reasoning and visual question answering (VQA) tasks.
DocVQA (Document Visual Question Answering) Mathew et al. (2021) DocVQA focuses on visual question answering over document images, where questions pertain to scanned documents, OCR-recognized text, and textual reasoning. This dataset is valuable for document comprehension, text recognition, and multimodal reasoning.
DVQA (Diagrammatic Visual Question Answering) Kafle et al. (2018) DVQA is designed for visual question answering over diagrams and charts, covering questions related to bar charts, pie charts, and scientific illustrations. It evaluates the model’s ability to read structured visual information and perform reasoning based on graphical representations.
CLEVR (Compositional Language and Elementary Visual Reasoning) Johnson et al. (2017) CLEVR is a synthetic dataset for visual reasoning, containing 3D-rendered scenes with structured questions that require reasoning based on attributes, object relationships, and compositional logic. It is widely used to assess a model’s capability in compositional and multi-step reasoning.
TallyQA Acharya et al. (2019) TallyQA is a dataset specifically designed for object counting tasks, where questions require the model to accurately count objects in an image. It evaluates the model’s ability to attend to relevant objects, integrate global and local information, and perform numerical reasoning.
B.2 Evaluation Metrics